# The Origin Symphony: Probing Baryogenesis with Gravitational Waves

### v1

@@ I. AD scalars: background dynamics and asymmetry generation
In some versions of the Affleck-Dine mechanism, an extra coupling is introduced in the potential in order to generate a radial minimum at large field values Allahverdi and Mazumdar (2012). We argue that this is not necessary for the AD field to acquire a large initial value. It has been known that a light scalar field in de-Sitter space-time acquires quantum fluctuations, eventually leading to an equilibrium distribution Starobinsky and Yokoyama (1994). In the case of a quartic real scalar field with $V={1\over 4}\lambda\phi^{4}$, the variance of the equilibrium distribution is $\langle\phi^{2}\rangle\sim H^{2}/\sqrt{\lambda}$, whereas for a quadratic field $V={1\over 2}m^{2}\phi^{2}$ we arrive at $\langle\phi^{2}\rangle\sim H^{4}/m^{2}$. As an example, $\langle\phi^{2}\rangle\sim M_{\mathrm{Pl}}^{2}$ and $0.1\,\mathrm{GeV}\lesssim m\lesssim 10\,\mathrm{GeV}$ requires an inflationary Hubble scale of at least $10^{-10}M_{\mathrm{Pl}}\lesssim H\lesssim 10^{-9}M_{\mathrm{Pl}}$ for quadratic fields. If the quartic term dominates during inflation with $\lambda\sim 10^{-30}$ a Planckian standard deviation requires $H\sim 10^{-15}M_{\mathrm{Pl}}$ or greater. The variance of the distribution grows as $\langle\phi^{2}\rangle=(H^{2}/4\pi^{2})N$ where $N$ is the $e$-folding number. We thus see that an equilibrium distribution with a Planckian standard deviation requires much more than the usual minimum amount of $60$ $e$-folds of inflation. This may be considered as a way to infer whether a prolonged inflationary phase took place in the early universe.
The above discussion (see Ref. Cui et al. (2024) for more details) applies to the absolute value of the complex scalar field $|\phi|$. The argument will enter its equilibrium distribution much sooner, and the distribution itself will be much simpler: uniform between 00 and $2\pi$. The intricacies of multi-dimensional random walks and their applications to de Sitter fluctuations of fields with continuous symmetries (including complex scalars) can be found in Ref. Adshead et al. (2020). Finally, a large initial field value $\phi_{\mathrm{in}}$ could simply result from an ad hoc initial condition. One may argue that the initial condition for the inflaton itself is set ad hoc, although quantum fluctuations can be invoked to explain it. Most importantly, our results are independent of the details of how $\phi_{\mathrm{in}}$ is realized (the same is true for the inflationary attractor).
@@
We follow previous computations of isocurvature fluctuations in models of AD baryogenesis, in particular, Ref. Hertzberg and Karouby (2014); see also Kasuya et al. (2008). During inflation we can decompose the AD field $\Phi$ into a radial component $R$ and an angular component $\theta$. We consider a typical value of the radial field $\bar{R}=\sqrt{\langle R^{2}\rangle}\sim M_{\mathrm{P}l}$ and thus
@@
where we have used the subscript ‘00’ to denote the generated asymmetry in the $\Phi$ sector alone, without considering the decay to SM particles.
@@
This expression has two limiting cases depending on whether the $\Phi$ lifetime is short ($\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}$) or long ($\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}$) compared to the period of oscillation of the asymmetry:
@@
To this end, we multiply the comoving transferred asymmetry with $(a_{\ast}/a)^{3}$ to compute the total asymmetry
@@
For the RD universe, as in our case, the equation of state is $w=1/3$ leading to
@@
$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{k_{T_{d}}^{6}\lambda_{{}_{\Phi}}}\right)^{1/4}\left(\frac{A}{m_{{}_{\Phi}}^{2}}\right)\left(\frac{m_{{}_{\Phi}}M_{\mathrm{P}}}{T_{d}^{2}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{{}_{\Phi}}}\right)^{1/4}\left(\frac{m_{{}_{\Phi}}^{2}}{A}\right)\left(\frac{T_{d}^{2}}{m_{{}_{\Phi}}M_{\mathrm{P}}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}\,,\end{cases}$ (S26)
@@
$\frac{n_{B}}{s}=\begin{cases}\mu\left(\frac{4\lambda_{{}_{\Phi}}\alpha^{3}}{k_{T_{d}}^{6}}\right)^{1/4}\left(\frac{A}{m_{{}_{\Phi}}^{2}}\right)\left(\frac{\phi_{\mathrm{in}}M_{\mathrm{P}}}{T_{d}^{2}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}\\ \mu\left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{{}_{\Phi}}^{3}}\right)^{1/4}\left(\frac{m_{{}_{\Phi}}^{2}}{A}\right)\left(\frac{T_{d}^{2}}{\phi_{\mathrm{in}}M_{\mathrm{P}}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}\,,\end{cases}$ (S27)
@@
In the case where the asymmetry in the $\Phi$ field is first transferred to a SM lepton asymmetry $n_{L}$, which is later converted to a baryon asymmetry $n_{B}$ via the sphaleron process, the above result gets multiplied by a factor of $28/79$ since Harvey and Turner (1990); Asaka et al. (2002); Garcia-Bellido et al. (1999)
$\frac{n_{B}}{s}=-\frac{28}{79}\frac{n_{L}}{s}\,.$ (S28)
Figure S1 shows the resulting asymmetry as a function of the temperature of the universe at the time of complete $\Phi$ decay for different values of the potential parameters; $m_{\Phi}$ and $A$. The quartic term is chosen to keep $m_{\Phi}^{2}/\lambda_{\Phi}\phi_{\mathrm{in}}^{2}=10^{-5}$ for sufficient resonance. We see the strong dependence of $n_{B}/s$ on $A$ and $T_{d}$, as expected. This allows the model to be viable, even in cases where significant washout occurs, by increasing the initially produced asymmetry through proper parameter choices.
@@
Finally, let us reiterate one assumption that was made in the above analytic results for the generated asymmetry. As mentioned in the main text, our analysis assumes that the motion of $\Phi$ is radial at early times, for $\lambda_{\Phi}|\Phi|^{4}\gg m_{\Phi}^{2}|\Phi|^{2}$. After that, the potential is dominated by the quadratic terms, which in our case includes the bare mass and the asymmetry $A(\Phi^{2}+\Phi_{*}^{2})$. By simply equating the quartic and bare mass terms, we arrive at the time of transition between the two regimes of $t_{*}$ when $|\Phi(t_{*})|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$. Demanding that at early times (before $t_{*}$) the asymmetry term can be neglected leads to $\epsilon_{\Phi}=A/m_{\phi}^{2}<(\mu^{2}/15\alpha)(\phi_{\mathrm{in}}^{2}/M_{\mathrm{Pl}}^{2})$ where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})\lesssim 10^{-5}$. For $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$ the above inequality can be written as $\epsilon_{\Phi}\lesssim 10^{-5}$. If we push $\epsilon_{\Phi}$ above this value, the model is not invalided, but the asymmetry would get an extra suppression factor since $\sin\theta<\sin\theta_{\mathrm{in}}$. The amount of suppression depends on the amount of non-radial motion for $t<t_{*}$. That being said, the computed $n_{B}/s$ can be many orders of magnitude above the measured value of $10^{-10}$, meaning that a suppression of $\sin\theta$ is not “fatal” for the model, even for $\epsilon_{\Phi}>10^{-5}$. We do not perform a detailed analysis for this regime, which simply involves computing the angular motion of $\Phi$ from the first instant when it starts rolling for a given set of potential parameters.
@@
As an example, for the benchmark values of $\lambda_{\Phi}\sim 10^{-30}$ and $m_{\Phi}=10\,\mathrm{GeV}$, we must have $\epsilon_{\Phi}\lesssim 10^{-6}$ for the validity of our estimates in Eq. (5). The temperature at the decay can vary between $T_{d}^{\mathrm{max}}\sim 10^{8}\,\mathrm{GeV}$ for instantaneous decay and $T_{d}^{\mathrm{min}}=10\mathrm{MeV}$, if the decay occurs right before BBN. Therefore, to obtain the correct ratio of $n_{B}/s$, we require $\epsilon_{\Phi}=10^{-38}(T_{d}/\mathrm{GeV})^{2}$ when $\gamma_{\Phi}\ll\epsilon_{\Phi}$.
@@
Figure S1: The maximum baryon-to-photon ratio $\left.{n_{B}/s}\right|_{\mathrm{max}}={n_{\Phi}/s}$ (neglecting possible washout and the sphaleron factor) for three different masses of the scalar field as in Fig (1) as a function of the temperature $T_{d}$ of the $\Phi$ decay. The solid and dashed lines correspond to $\epsilon_{\Phi}=10^{-20}$ and $\epsilon_{\Phi}=10^{-6}$ respectively. The maximum temperature (denoted as the dots) for each line corresponds to the case of instantaneous decay of the scalar at $t=t_{\ast}$. The horizontal dot-dashed line corresponds to the observed value of the baryon asymmetry.
@@
In this section we elaborate the example models realizing the transfer of $\Phi$ asymmetry to the SM baryon asymmetry and the related constraints.
@@
We consider the simple scenario where $\Phi$ decays to $NN$, then $N$ transfers the asymmetry to SM leptons via freeze-out or freeze-in processes as considered in leptogenesis literature Akhmedov et al. (1998); Klarić et al. (2021a, b); Cui et al. (2011); Flood et al. (2022); Shuve and Yavin (2014); Hambye and Teresi (2016). As seen from our GW results, $m_{\Phi}\lesssim O(10)$ GeV is preferred in order to yield detectable GW frequency below $\sim 100$ Hz. This implies that light $N$ of mass $\lesssim O(1-10)$ GeV. In this mass range, conventional $N$ decays to SM leptons via a 3-body or loop-suppressed process, which typically occurs after electroweak sphaleron already turns off, albeit before BBN. The washout effect from inverse decay or back-scattering is also significant in this low mass range. However, as found in recent leptogenesis literature, asymmetry transfer from $N$ to SM leptons can be realized while sphaleron is still active by freeze-out (generalized out-of-equilibrium decay or scattering) or freeze-in (via $N$-$\nu_{\mathrm{S}M}$ mixing/oscillation) mechanisms, for $m_{N}$ as low as O(100) MeV. The strength of the washout effect can be highly flavor-dependent. Even in the presence of large washout suppression, as shown in the main text, our Affleck-Dine model can initially generate a larger $\Phi$ asymmetry, which compensates for the washout to eventually result in the observed baryon asymmetry. Unlike with the leptogeneis case, here, the $NHL$ interaction only serves the purpose of transferring the $\Phi$ ($N$) asymmetry to SM leptons instead of generating the asymmetry via CP violation in the SM neutrino sector. Therefore, the details of the asymmetry evolution in our case may differ from its leptogenesis counterparts and are highly model-dependent, which may be elucidated by dedicated studies beyond the scope of this work. Furthermore, an alternative to the Majorana $N$ as shown in Eq. S29 is to introduce quasi-Dirac $N$ where the washout effect can be more suppressed for GeV mass range (also see Cline et al. (2020b) for this possibility and the related phenomenological implications). For all of the above possibilities for $N$ to SM asymmetry transfer, $\Phi\rightarrow NN$ decay is required to occur before EW sphaleron turns off, i.e. $\Gamma_{\Phi\rightarrow NN}\sim y_{\Phi}^{2}m_{\Phi}/(16\pi)\gtrsim H(T_{\mathrm{E}W})$. This leads to the constraint $y_{\Phi}\gtrsim 10^{-7}\sqrt{\frac{\mathrm{GeV}}{m_{\Phi}}}$, which can be easily satisfied.
Laboratory signatures of the case of $\Phi$ decaying to leptons generally relate to sterile neutrino searches in the mass range of $O(0.1-10)$ GeV, which are being conducted at a range of experiments such as DUNE, SHiP, FASER et al. ste (2022); Bolton et al. (2020); Ariga et al. (2019); Ballett et al. (2020); Abe et al. (2019); Drewes et al. (2018); Barouki et al. (2022).
@@
where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices $i,~{}j=1,2,3$ in down-type quark combination $d_{i}d_{j}$ should be antisymmetric. This type of model finds its close analogy in supersymmetric theories, in particular, $\chi$ may relate to neutralino or singlino Cui and Sundrum (2013), $\varphi$ an up-type squark, $\lambda_{ij}$ as UDD type of R-parity violating (RPV) coupling. Therefore, existing searches for RPV SUSY may lead to constraints on masses and couplings that we consider here. The LHC constraints on the colored particle $\phi$ depend on whether it resembles the first two generations of squark or stop, and depends on whether RPV decay $\phi\rightarrow dd$ or R-parity conserving decay $\phi\rightarrow u\chi$ dominates (assuming $\chi$ reveals itself as MET) Karmakar (2023); Sirunyan et al. (2018, 2021); Aad et al. (2021). In any case, generally, $m_{\varphi}\gtrsim$1 TeV is the ballpark of the current LHC constraints, which we will use for our estimates. This constraint determines that in the parameter range of our interest where $m_{\chi}\sim O(1-10)$ GeV, $\chi$ decays proceeds with 3-body channel $\chi\rightarrow udd$. For simplicity, we assume $g_{i}\sim 1$. B-violating coupling $\lambda_{ij}$ is subject to various constraints by laboratory searches at the intensity frontier. This model does not involve L-violation and does not induce proton decay, which usually introduces constraints on B-violating couplings Arnold et al. (2013). Due to the antisymmetry in $i,~{}j$, $n$-$\bar{n}$ constraint is rather weak, while di-nucleon decay $pp\rightarrow K^{+}K^{+}$ provides the strongest bound Goity and Sher (1995); Barbier et al. (2005) of $\lambda_{12}\lesssim 10^{-6}$, for $m_{\varphi}\sim$ TeV, $m_{\chi}\sim$ O(GeV). However, the bound is relaxed for heavy flavor quarks, e.g. $\lambda_{23}\lesssim 1$.
@@
The kinematic condition of $\chi$ decaying to baryons leads to lower limit on masses: $m_{\Phi}>2$ GeV, $m_{\chi}>1$ GeV. The decays of both $\Phi\rightarrow\chi\chi$ and $\chi\rightarrow udd$ are required to occur before BBN, i.e., $\Gamma_{\Phi\rightarrow\chi\chi}\sim\frac{1}{16\pi}y_{\chi}^{2}m_{\Phi}>H(T_{\mathrm{BBN}})$, and $\Gamma_{\chi\rightarrow udd}>H(T_{\mathrm{BBN}})$. We then estimate the constraint on coupling parameters as: $y_{\chi}\gtrsim 10^{-12}\frac{\mathrm{G}eV}{m_{\phi}}$, $\sum_{i,j}g_{i}\lambda_{ij}\gtrsim\sqrt{10^{-21}\frac{m^{4}_{\varphi}\mathrm{GeV}}{m_{\chi}^{5}}}$. With $m_{\varphi}\sim$ TeV, $m_{\chi}\sim$ O(GeV), we find that $\sum_{i,j}g_{i}\lambda_{ij}\gtrsim 10^{-5}$. If $\lambda_{12}$ dominates $\chi$ decay, this would contradict the aforementioned di-nucleon decay bound. Therefore, we choose to consider the pattern of heavy flavor domination, which is well motivated and has been widely considered in SUSY context Brust et al. (2012a, b); Cui et al. (2011); Franceschini and Torre (2013). In particular, we consider the dominant decay of $\chi$ involves b-quark $\chi\rightarrow udb$ or $\chi\rightarrow usb$. As $m_{b}\simeq 4$ GeV, the kinematics requires $m_{\chi}>$4 GeV and $m_{\Phi}>$8 GeV. The involvement of heavy b-quark also alleviates the potential washout suppression of baryon asymmetry due to inverse decay $udd\rightarrow\chi$ and back-scattering $udd\rightarrow\bar{u}\bar{d}\bar{d}$: by requiring $\chi$ decay below $T\sim m_{b}$ (compatible with the condition of decay before BBN), the washout effect gets Boltzmann suppression and quickly becomes inefficient.

### v2

@@ Abstract
In some versions of the Affleck-Dine mechanism, an extra coupling is introduced in the potential in order to generate a radial minimum at large field values Allahverdi and Mazumdar (2012). We argue that this is not necessary for the AD field to acquire a large initial value. It has been known that a light scalar field in de-Sitter space-time acquires quantum fluctuations, eventually leading to an equilibrium distribution Starobinsky and Yokoyama (1994). In the case of a quartic real scalar field with $V={1\over 4}\lambda\phi^{4}$, the variance of the equilibrium distribution is $\langle\phi^{2}\rangle\sim H^{2}/\sqrt{\lambda}$, whereas for a quadratic field $V={1\over 2}m^{2}\phi^{2}$ we arrive at $\langle\phi^{2}\rangle\sim H^{4}/m^{2}$. As an example, $\langle\phi^{2}\rangle\sim M_{\mathrm{Pl}}^{2}$ and $0.1\,\mathrm{GeV}\lesssim m\lesssim 10\,\mathrm{GeV}$ requires an inflationary Hubble scale of at least $10^{-10}M_{\mathrm{Pl}}\lesssim H\lesssim 10^{-9}M_{\mathrm{Pl}}$ for quadratic fields. If the quartic term dominates during inflation with $\lambda\sim 10^{-30}$ a Planckian standard deviation requires $H\sim 10^{-15}M_{\mathrm{Pl}}$ or greater. The variance of the distribution grows as $\langle\phi^{2}\rangle=(H^{2}/4\pi^{2})N$ where $N$ is the $e$-folding number. We thus see that an equilibrium distribution with a Planckian standard deviation requires much more than the usual minimum amount of 6060 $e$-folds of inflation. This may be considered as a way to infer whether a prolonged inflationary phase took place in the early Universe.
The above discussion (see Ref. Cui et al. (2024) for more details) applies to the absolute value of the complex scalar field $|\phi|$. The argument will enter its equilibrium distribution much sooner, and the distribution itself will be much simpler: uniform between 0 and $2\pi$. The intricacies of multi-dimensional random walks and their applications to de Sitter fluctuations of fields with continuous symmetries (including complex scalars) can be found in Ref. Adshead et al. (2020). Finally, a large initial field value $\phi_{\mathrm{in}}$ could simply result from an ad hoc initial condition. One may argue that the initial condition for the inflaton itself is set ad hoc, although quantum fluctuations can be invoked to explain it. Most importantly, our results are independent of the details of how $\phi_{\mathrm{in}}$ is realized (the same is true for the inflationary attractor).
@@
We follow previous computations of isocurvature fluctuations in models of AD baryogenesis, in particular, Ref. Hertzberg and Karouby (2014); see also Kasuya et al. (2008). During inflation, we can decompose the AD field $\Phi$ into a radial component $R$ and an angular component $\theta$. We consider a typical value of the radial field $\bar{R}=\sqrt{\langle R^{2}\rangle}\sim M_{\mathrm{P}l}$ and thus
@@
where we have used the subscript ‘0’ to denote the generated asymmetry in the $\Phi$ sector alone, without considering the decay to SM particles.
@@
This expression has two limiting cases, depending on whether the $\Phi$ lifetime is short ($\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}$) or long ($\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}$) compared to the period of oscillation of the asymmetry:
@@
To this end, we multiply the comoving transferred asymmetry by $(a_{\ast}/a)^{3}$ to compute the total asymmetry
@@
For the RD Universe, as in our case, the equation of state is $w=1/3$, leading to
@@
[…]
@@
[…]
@@
When written in terms of the dimensionless parameters, $\epsilon_{\Phi}=A/m_{\Phi}^{2}$ and $\gamma_{\Phi}=\Gamma_{\Phi}/m_{\Phi}$, we find the following familiar expression:
$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{\lambda_{\Phi}k_{T_{d}}^{6}}\right)^{1/4}\epsilon_{\Phi}\frac{m_{\Phi}M_{\mathrm{Pl}}}{T_{d}^{2}}\sin(2\theta);&\gamma_{\Phi}\gg 2\epsilon_{\Phi}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{\Phi}}\right)^{1/4}\frac{1}{\epsilon_{\Phi}}\frac{T_{d}^{2}}{m_{\Phi}M_{\mathrm{Pl}}}\sin(2\theta);&\gamma_{\Phi}\ll 2\epsilon_{\Phi}\end{cases}$ (S28)
@@
Figure S1 shows the resulting asymmetry, following Eq. (S22), as a function of the temperature of the Universe at the time of complete $\Phi$ decay for different values of the potential parameters; $m_{\Phi}$ and $A$. The quartic term is chosen to keep $m_{\Phi}^{2}/\lambda_{\Phi}\phi_{\mathrm{in}}^{2}=10^{-5}$ for sufficient resonance. We see the strong dependence of $n_{B}/s$ on $A$ and $T_{d}$, as expected. This enables the model to remain viable, even in cases where significant washout occurs, by increasing the initially produced asymmetry through appropriate parameter choices. Finally, let us reiterate one assumption made in the above analytic results regarding the generated asymmetry. As mentioned in the main text, our analysis assumes that the motion of $\Phi$ is radial at early times, for $\lambda_{\Phi}|\Phi|^{4}\gg m_{\Phi}^{2}|\Phi|^{2}$. After that, the potential is dominated by the quadratic terms, which in our case include the bare mass and the asymmetry $A(\Phi^{2}+\Phi_{*}^{2})$. By simply equating the quartic and bare mass terms, we arrive at the time of transition between the two regimes of $t_{*}$ when $|\Phi(t_{*})|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$. Demanding that at early times (before $t_{*}$) the asymmetry term can be neglected leads to $\epsilon_{\Phi}<(\mu^{2}/15\alpha)(\phi_{\mathrm{in}}^{2}/M_{\mathrm{Pl}}^{2})$ where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})\lesssim 10^{-5}$. For $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$ the above inequality can be written as $\epsilon_{\Phi}\lesssim 10^{-5}$. If we push $\epsilon_{\Phi}$ above this value, the model is not invalidated, but the asymmetry would get an extra suppression factor since $\sin\theta<\sin\theta_{\mathrm{in}}$. The amount of suppression depends on the amount of non-radial motion for $t<t_{*}$. That being said, the computed $n_{B}/s$ can be many orders of magnitude above the measured value of $10^{-10}$, meaning that a suppression of $\sin\theta$ is not “fatal” for the model, even for $\epsilon_{\Phi}>10^{-5}$. We do not perform a detailed analysis for this regime, which involves simply computing the angular motion of $\Phi$ from the first instant when it starts rolling, given a specific set of potential parameters.
@@
As an example, for the benchmark values of $\lambda_{\Phi}\sim 10^{-30}$ and $m_{\Phi}=10\,\mathrm{GeV}$, we must have $\epsilon_{\Phi}\lesssim 10^{-6}$ for the validity of our estimates in Eq. (S27). The temperature at the decay can vary between $T_{d}^{\mathrm{max}}\sim 10^{8}\,\mathrm{GeV}$ for instantaneous decay and $T_{d}^{\mathrm{min}}=10\mathrm{MeV}$, if the decay occurs right before BBN Kawasaki et al. (1999, 2000); Steigman (2007); Hasegawa et al. (2019). Therefore, to obtain the correct ratio of $n_{B}/s$, we require $\epsilon_{\Phi}=10^{-38}(T_{d}/\mathrm{GeV})^{2}$ when $\gamma_{\Phi}\ll\epsilon_{\Phi}$.
@@
Figure S1: The maximum baryon-to-photon ratio $\left.{n_{B}/s}\right|_{\mathrm{max}}={n_{\Phi}/s}$ (neglecting possible washout and the sphaleron factor) for three different masses of the scalar field, as in Fig. (1) of the main text, as a function of the temperature $T_{d}$ of the $\Phi$ decay. The solid and dashed lines correspond to $\epsilon_{\Phi}=10^{-20}$ and $\epsilon_{\Phi}=10^{-6}$ respectively. The maximum temperature (denoted as the dots) for each line corresponds to the case of instantaneous decay of the scalar at $t=t_{\ast}$. The horizontal dot-dashed line corresponds to the observed value of the baryon asymmetry.
@@
In this section, we elaborate on the example models realizing the transfer of $\Phi$ asymmetry to the SM baryon asymmetry and the related constraints.
@@
We consider the simple scenario where $\Phi$ decays to $NN$, then $N$ transfers the asymmetry to SM leptons via freeze-out or freeze-in processes as considered in leptogenesis literature Akhmedov et al. (1998); Klarić et al. (2021a, b); Cui et al. (2011); Flood et al. (2022); Shuve and Yavin (2014); Hambye and Teresi (2016). As seen from our GW results, $m_{\Phi}\lesssim O(10)$ GeV is preferred in order to yield a detectable GW frequency below $\sim 100$ Hz. This implies that light $N$ of mass $\lesssim O(1-10)$ GeV. In this mass range, conventional $N$ decays to SM leptons via a 3-body or loop-suppressed process, which typically occurs after electroweak sphaleron already turns off, albeit before BBN. The washout effect from inverse decay or backscattering is also significant in this low-mass range. However, as found in recent leptogenesis literature, asymmetry transfer from $N$ to SM leptons can be realized while sphaleron is still active by freeze-out (generalized out-of-equilibrium decay or scattering) or freeze-in (via $N$-$\nu_{\mathrm{S}M}$ mixing/oscillation) mechanisms, for $m_{N}$ as low as O(100) MeV. The strength of the washout effect can be highly flavor-dependent. Even in the presence of large washout suppression, as shown in the main text, our Affleck-Dine model can initially generate a larger $\Phi$ asymmetry, which compensates for the washout to eventually result in the observed baryon asymmetry. Unlike with the leptogenesis case, here, the $NHL$ interaction only serves the purpose of transferring the $\Phi$ (NN) asymmetry to SM leptons instead of generating the asymmetry via CP violation in the SM neutrino sector. Therefore, the details of the asymmetry evolution in our case may differ from those of its leptogenesis counterparts and are highly model-dependent, which can be elucidated by dedicated studies beyond the scope of this work. Furthermore, an alternative to the Majorana $N$ as shown in Eq. S29 is to introduce quasi-Dirac $N$, where the washout effect can be more suppressed for the GeV mass range (also see Cline et al. (2020b) for this possibility and the related phenomenological implications). For all of the above possibilities for $N$ to SM asymmetry transfer, $\Phi\rightarrow NN$ decay is required to occur before EW sphaleron turns off, i.e. $\Gamma_{\Phi\rightarrow NN}\sim y_{\Phi}^{2}m_{\Phi}/(16\pi)\gtrsim H(T_{\mathrm{E}W})$. This leads to the constraint $y_{\Phi}\gtrsim 10^{-7}\sqrt{\frac{\mathrm{GeV}}{m_{\Phi}}}$, which can be easily satisfied.
Laboratory signatures of the case of $\Phi$ decaying to leptons generally relate to sterile neutrino searches in the mass range of O(0.110)O(0.1-10) GeV, which are being conducted at a range of experiments such as DUNE, SHiP, FASER et al. ste (2022); Bolton et al. (2020); Ariga et al. (2019); Ballett et al. (2020); Abe et al. (2019); Drewes et al. (2018); Barouki et al. (2022).
@@
where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices i,j=1,2,3i,~j=1,2,3 in down-type quark combination $d_{i}d_{j}$ should be antisymmetric. This type of model finds its close analogy in supersymmetric theories. In particular, $\chi$ may relate to a neutralino or singlino Cui and Sundrum (2013), $\varphi$ to an up-type squark, and $\lambda_{ij}$ to a UDD-type R-parity-violating (RPV) coupling. Therefore, existing searches for RPV SUSY may lead to constraints on masses and couplings that we consider in this context. The LHC constraints on the colored particle $\phi$ depend on whether it resembles the first two generations of squark or stop, and depend on whether RPV decay $\phi\rightarrow dd$ or R-parity conserving decay $\phi\rightarrow u\chi$ dominates (assuming $\chi$ reveals itself as MET) Karmakar (2023); Sirunyan et al. (2018, 2021); Aad et al. (2021). In any case, generally, $m_{\varphi}\gtrsim$1 TeV is the ballpark of the current LHC constraints, which we will use for our estimates. This constraint determines that in the parameter range of our interest, where $m_{\chi}\sim O(1-10)$ GeV, $\chi$ decays proceed with the 3-body channel $\chi\rightarrow udd$. For simplicity, we assume $g_{i}\sim 1$. B-violating coupling $\lambda_{ij}$ is subject to various constraints by laboratory searches at the intensity frontier. This model does not involve L-violation and does not induce proton decay, which usually introduces constraints on B-violating couplings Arnold et al. (2013). Due to the antisymmetry in i,ji,~j, $n¯\bar{n}$ constraint is rather weak, while di-nucleon decay $pp\rightarrow K^{+}K^{+}$ provides the strongest bound Goity and Sher (1995); Barbier et al. (2005) of $\lambda_{12}\lesssim 10^{-6}$, for $m_{\varphi}\sim$ TeV, $m_{\chi}\sim$ O(GeV). However, the bound is relaxed for heavy flavor quarks, e.g., $\lambda_{23}\lesssim 1$.
@@
The kinematic condition of $\chi$ decaying to baryons leads to a lower limit on masses: $m_{\Phi}>2$ GeV, $m_{\chi}>1$ GeV. The decays of both $\Phi\rightarrow\chi\chi$ and $\chi\rightarrow udd$ are required to occur before BBN, i.e., $\Gamma_{\Phi\rightarrow\chi\chi}\sim\frac{1}{16\pi}y_{\chi}^{2}m_{\Phi}>H(T_{\mathrm{BBN}})$, and $\Gamma_{\chi\rightarrow udd}>H(T_{\mathrm{BBN}})$. We then estimate the constraint on coupling parameters as: $y_{\chi}\gtrsim 10^{-12}\frac{\mathrm{G}eV}{m_{\phi}}$, $\sum_{i,j}g_{i}\lambda_{ij}\gtrsim\sqrt{10^{-21}\frac{m^{4}_{\varphi}\mathrm{GeV}}{m_{\chi}^{5}}}$. With $m_{\varphi}\sim$ TeV, $m_{\chi}\sim$ O(GeV), we find that $\sum_{i,j}g_{i}\lambda_{ij}\gtrsim 10^{-5}$. If $\lambda_{12}$ dominates $\chi$ decay, this would contradict the aforementioned di-nucleon decay bound. Therefore, we choose to consider the pattern of heavy flavor domination, which is well-motivated and has been widely studied in the SUSY context Brust et al. (2012a, b); Cui et al. (2011); Franceschini and Torre (2013). In particular, we consider the dominant decay of $\chi$ involves b-quark $\chi\rightarrow udb$ or $\chi\rightarrow usb$. As $m_{b}\simeq 4$ GeV, the kinematics requires $m_{\chi}>$4 GeV and $m_{\Phi}>$8 GeV. The involvement of heavy b-quark also alleviates the potential washout suppression of baryon asymmetry due to inverse decay $udd\rightarrow\chi$ and back-scattering $udd\rightarrow\bar{u}\bar{d}\bar{d}$: by requiring $\chi$ decay below $T\sim m_{b}$ (compatible with the condition of decay before BBN), the washout effect gets Boltzmann suppression and quickly becomes inefficient.